OpenStudy (anonymous):

write sqr36 * 3sqr64 using exponential notation

OpenStudy (anonymous):

write \[\sqrt{36}\times \sqrt[3]{64}\] in exponential notation

OpenStudy (dan815):

sqrt of 36 =?

OpenStudy (dan815):

can you think of a number multiplied by itself that gives 6

OpenStudy (mathstudent55):

\( \Large x^{\frac{1}{a}} = \sqrt[a]{x } \)

OpenStudy (dan815):

36*

OpenStudy (dan815):